# Entropy-Lens: The Information Signature of Transformer Computations

### v2

@@ Abstract
Transformer models map input token sequences to output token distributions, layer by layer. While most interpretability work focuses on internal latent representations, we study the evolution of these token-level distributions directly in vocabulary space. However, such distributions are high-dimensional and defined on an unordered support, making common descriptors like moments or cumulants ill-suited. We address this by computing the Shannon entropy of each intermediate predicted distribution, yielding one interpretable scalar per layer. The resulting sequence—the entropy profile—serves as a compact, information-theoretic signature of the model’s computation.
We introduce Entropy-Lens, a model-agnostic framework that extracts entropy profiles from frozen, off-the-shelf transformers. We show that these profiles (i) reveal family-specific computation patterns invariant under depth rescaling, (ii) are predictive of prompt type and task format, and (iii) correlate with output correctness. We further show that Rényi entropies yield similar results within a broad range of $\alpha$ values, justifying the use of Shannon entropy as a stable and principled summary. Our results hold across different transformers, without requiring gradients, fine-tuning, or access to model internals.
00footnotetext: Preprint - This manuscript is currently under review.11footnotetext: Code available at: github.com/christopher-irw/Entropy-Lens
@@ 1 Introduction
Transformer-based architectures (Vaswani et al. 2023) operate by iteratively mapping token inputs to distributions over next-token outputs, layer by layer (Shan et al. 2024). This view emphasizes their inherently probabilistic nature: each layer produces a distribution over the vocabulary, encoding the model’s beliefs at that stage.
Despite their success across domains—from language to biology to vision (Devlin 2018; Ji et al. 2021; Wu et al. 2020)—the internal evolution of these distributions remains poorly understood, resulting in unpredictable behaviour (Wei et al. 2022) and reliability concerns (Schroeder and Wood-Doughty 2025; Huang et al. 2025). Most interpretability research either focuses on toy models (Elhage et al. 2021, 2022) or requires training a set of probes (nostalgebraist 2020; Belrose et al. 2023) or full models on ad-hoc tasks (Nanda et al. 2023), limiting scalability and generality. Moreover, existing methods primarily study latent representations—hidden states, attention maps, or MLP activations (Nanda and Bloom 2022; Bereska and Gavves 2024; Chefer, Gur, and Wolf 2021)—rather than the output distributions themselves.
Figure 1: TSNE of the aggregated entropy profiles of different model families using unprompted models. The size of each point is scaled based on the number of parameters of the model.
We propose to shift perspective: instead of analyzing the internal geometry of latent activations, we focus on the evolution of the model’s token-level predictions across layers. These predictions, obtained by decoding intermediate representations via the model’s output head, form probability distributions over the vocabulary. Studying their evolution offers a direct, vocabulary-grounded view of the computation.
However, these distributions pose two challenges: they are high-dimensional (one value per token) and defined over an unordered support (the vocabulary). This makes classical descriptors like variance or higher-order cumulants unstable or ill-defined. We address both issues by computing the Shannon entropy of each distribution—a scalar, interpretable quantity invariant to token permutation, and reflective of the model’s uncertainty at each layer. We introduce Entropy-Lens, a simple, scalable, and model-agnostic framework for studying transformer computations through the evolution of entropy across layers.
In our experiments, we consider several LLMs, including Llama (Touvron et al. 2023), Gemma (Team et al. 2024), and GPT, (Radford 2018) up to 9B parameters. For each generated token, we compute the entropy of its intermediate predictions, as described in Section 4, yelding one interpretable value per layer: this constitutes the entropy profile. We then aggregate these profiles across tokens and study whether they capture structural information about the internal LLM’s computation (Figure 1). In Section 5, we show that: (i) entropy profiles reveal family-specific signatures of transformer architectures, invariant under depth rescaling; (ii) they are predictive of prompt semantics and format; (iii) they correlate with output correctness; (iv) and results are stable across Rényi entropy variants, justifying our use of Shannon entropy as a principled default.
Entropy-Lens requires no gradients, no fine-tuning, and no access to internal weights. It applies to frozen, off-the-shelf transformers of arbitrary size. While our experiments focus on language models, the methodology is general and opens new directions for vocabulary-grounded interpretability.

### v3

# Entropy-Lens: Uncovering Decision Strategies in LLMs

@@ Abstract
In large language models (LLMs), each block operates on the residual stream to map input token sequences to output token distributions. However, most of the interpretability literature focuses on internal latent representations, leaving token-space dynamics underexplored. The high dimensionality and categoricity of token distributions hinder their analysis, as standard statistical descriptors are not suitable. We show that the entropy of logit-lens predictions overcomes these issues. In doing so, it provides a per-layer scalar, permutation-invariant metric. We introduce Entropy-Lens to distill the token-space dynamics of the residual stream into a low-dimensional signal. We call this signal the entropy profile.
We apply our method to a variety of model sizes and families, showing that (i) entropy profiles uncover token prediction dynamics driven by expansion and pruning strategies; (ii) these dynamics are family-specific and invariant under depth rescaling; (iii) they are characteristic of task type and output format; (iv) these strategies have unequal impact on downstream performance, with the expansion strategy usually being more critical. Ultimately, our findings further enhance our understanding of the residual stream, enabling a granular assessment of how information is processed across model depth.
Machine Learning, ICML
@@ 1 Introduction
Transformer-based architectures (Vaswani et al., 2023) consist of stacked blocks. Each one operates on the residual stream by iteratively mapping input tokens to output distributions over the next-token (Shan et al., 2024). Much of the interpretability literature studies the residual stream in embedding space (Skean et al., 2025; Jawahar et al., 2019; Elhage et al., 2022), often by analyzing geometric arrangements and dynamics of aggregated latent representations (Skean et al., 2025; Sarfati et al., 2025; Li et al., 2025). These methods offer useful diagnostics but fail to offer a low-dimensional, interpretable, and per-sample representation of the model’s behavior in token space (Sarfati et al., 2025). We tackle this gap by projecting the residual stream into the token space via logit-lens (nostalgebraist, 2020) and analyzing the entropy dynamics of the decoded distributions. With this approach, grounded in information theory, we distill the residual stream into a low-dimensional signal. This allows us to track token prediction dynamics layer-by-layer and on individual generations.
Our framework, Entropy-Lens, uses entropy as a natural metric to uncover token prediction dynamics driven by two fundamental strategies: (1) expansion, where the model considers new candidates for the next token, and (2) pruning, where it refines the candidate set. We compute the entropy of the logit-lens probabilities as a count of the number of possible candidates for each layer and sample. Our analysis reveals that model families have a characteristic way of mixing these two strategies across layers, which is consistent between models of different sizes in the same family. Furthermore, we show that different tasks (e.g. generative, syntactic, semantic) and output-formats (e.g. poem, scientific piece, chat log) induce distinct token prediction dynamics. Finally, we show in a multiple-choice setup that expansion and pruning strategies have different relative importance for downstream performance. At the same time, we observe an overall trend emphasizing the expansion phases.
In summary, our main contributions are:
  • We propose entropy as a sound metric to study the token prediction dynamics and expansion/pruning strategies in the transformer body with single layer and single token level granularity.
  • We show that the particular mixing between these strategies is characteristic of model families, task type and output format.
  • We show that, given a particular model, expansion/pruning strategies have different relative importance for downstream performance, with an overall trend emphasizing expansion.
2 Background
In this section, we review the fundamental concepts from information theory and key components of the Transformer architecture, with particular emphasis on the residual connections and the information flow across layers. We also briefly categorize the training stages of Large Language Models to clarify the distinction between the base and instruct models used in our study.
2.1 Information theory
Shannon entropy. The main information-theoretic quantity used in our study is entropy. Given a discrete [^1] Although entropy can be naturally extended to the continuous case with probability density functions, we restrict ourselves to the discrete case as it is the most relevant to our study. random variable $X$ with outcomes $x_{i}$ and probability mass function $p$, the Shannon entropy $H$ of $X$ is defined as
$H(X)=-\sum_{i}p(x_{i})\log p(x_{i})=\mathbb{E}[-\log p(X)].$ (1)
Shannon proved that this function is the only one, up to a scalar multiplication, which satisfies intuitive properties for measuring ‘disorder’ (Shannon, 1948). These include being maximal for a uniform distribution, minimal for the limit of a Kronecker delta function, and ensuring that $H(A,B)\leq H(A)+H(B)$ for every possible random variable $A$ and $B$. The same function already existed in continuous form in physics, where it linked the probabilistic formalism of statistical mechanics with the more phenomenological framework of thermodynamics, where the term ‘entropy’ was originally coined (Gibbs, 1902).
Rényi entropy. In addition to Shannon entropy, we also consider its generalization known as Rényi entropy. Given a discrete random variable $X$ with probability mass function $p$, the Rényi entropy of order $\alpha>0$, $\alpha\neq 1$, is defined as:
$H_{\alpha}(X)=\frac{1}{1-\alpha}\log\sum_{i}p(x_{i})^{\alpha}.$ (2)
This formulation reduces to Shannon entropy in the limit $\alpha\to 1$, and introduces a tunable parameter $\alpha$ that modulates the sensitivity of the entropy to the distribution’s tail. Rényi entropy also subsumes many classical descriptors of discrete distributions without intrinsic ordering: with appropriate choices of $\alpha$, it recovers collision entropy ($\alpha=2$), min-entropy ($\alpha\to\infty$), and max-entropy ($\alpha\to 0$), and it correlates with indices such as the Gini–Simpson index (Rényi, 1961; Jost, 2006). In general, lower values of $\alpha$ give more weight to rare events, while higher values emphasize the most probable outcomes.
Top-pp. Also referred to as Nucleus Sampling (Holtzman et al., 2020), Top-$p$ is a thresholding technique that restricts the distribution to the smallest set of tokens whose cumulative probability mass exceeds a value $p\in(0,1]$.
2.2 The Transformer
Architecture.
The transformer (Vaswani et al., 2023) is a deep learning architecture widely applied in language modelling with LLMs (Brown et al., 2020) and computer vision (Dosovitskiy et al., 2021). Transformer computations happen through transformer blocks and residual connections, as exemplified in Figure 2. While various design choices are possible, blocks are usually a composition of layer normalization (Zhang and Sennrich, 2019), attention, and multi layer perceptrons (MLPs). Residual connections, instead, sum the output of layer i1i-1 to the output of layer $i$.
Inside a single transformer block, the information flows both horizontally and vertically. The former, enabled by the attention mechanism, allows the token representations to interact with each other. In a language modelling task, for example, this is useful to identify which parts of the input sequence, the sentence prompt, should influence the next token prediction and quantify by how much. The latter vertical information flow allows the representation to evolve and encode different meanings or concepts. Usually, the dimension of the latent space is the same for each block in the transformer. The embedding spaces where these computations take place are generally called the residual stream.
Computation schema.
LLMs are trained to predict the next token in a sentence. That is, given a sentence prompt $S$ with tokens $t_{1},\dots,t_{N}$, the transformer encodes each token with a linear encoder $E$. Throughout the residual stream, the representation $\mathbf{x}_{N}$ of the token $t_{N}$ evolves into the representation of the token $t_{N+1}$, which is then decoded back into token space via a linear decoder $D$, often set to $E^{\top}$, tying the embedding weights and the decoder. Finally, the logits, the output of $D$, are normalized with $\operatorname{softmax}$ to represent a probability distribution over the vocabulary.
Instruct models. Training an LLM typically progresses from self-supervised pretraining on vast corpora to fine-tuning on structured instructions, often followed by RLHF for alignment (Ouyang et al., 2022). While distinction varies, Chat tuning focuses on dialogue history, whereas Instruct tuning (Wei et al., 2022) targets adherence to specific commands. In this work, we utilize off-the-shelf Instruct models (denoted ‘it’). We prioritize these over specific chat variants as their flexibility in following command-style prompts aligns better with our experimental setup.
3 Entropy and decision strategies
In this section, we motivate the use of entropy from two perspectives. First, Section 3.1 establishes it as a robust dimensionality reduction technique for high-dimensional, unordered vocabulary distributions. This addresses the main challenges posed by high-dimensional unordered distributions over token vocabularies. Second, Section 3.2 argue that changes in entropy $\Delta H_{i}$ between a layer $i$ and its predecessor provide a mechanistic interpretation, serving as indicators of expansion and pruning of the next-token candidate set. Finally, Section 3.3 generalizes our framework to the broader family of Rényi entropies. In particular, we argue that our findings are robust across a wide and informative range of $\alpha$ values, and not a simple arbitrary choice.
3.1 Entropy as dimensionality reduction
Analyzing distributions produced by transformers poses two challenges. First, they are high-dimensional, with one probability per vocabulary token. Second, there is no intrinsic notion of order between tokens in the vocabulary. As a result, many standard statistical summaries such as moments or cumulants are unstable or ill-defined when applied to vocabulary distributions.
We address both issues by computing the entropy of each distribution, which has two properties: (i) it is a scalar quantity, and (ii) it is invariant to token permutation. Property (i) addresses the high-dimensionality challenge, and property (ii) makes it suitable for unordered categorical distributions. In the main experiments, we use Shannon entropy $H$, defined in Section 2. Section 3.3 discusses the broader class of Rényi entropies satisfying the same properties, arguing that Shannon entropy can be taken as a robust parameter-free default.
3.2 Entropy changes as a proxy for expansion and pruning strategies
By definition, entropy is maximized when the probability mass is equally distributed, and minimized when it is concentrated in one event. In our setting, we focus on the entropy of layer-wise next-token predictions extracted via logit-lens (Section 4). Beyond its absolute value, we also analyze differences in the entropy of intermediate predictions between a layer $i$ and its predecessor $\Delta H_{i}=H_{i}-H_{i-1}$. We argue that we can interpret these increments and reductions as strategies of expansions and pruning of the candidate set, respectively. More formally, we define a candidate to be a token appearing in the top-$p$ distribution, with $p=0.6$. To this end, we need to validate two claims:
  1. C1
    $\Delta H_{i}$ is monotonically related to changes in the number of possible candidates: positive (negative) entropy variations correspond to increases (decreases) in the number of candidates, with larger variations indicating larger changes.
  2. C2
    Distributions across layers share a non-negligible portion of their candidates.
Claim C1 ensures that variations in entropy, $\Delta H$, reflect changes in the size of the top-$p$ candidate set, rather than fluctuations confined to the bottom-(1p)(1-p) tail of the distribution. Claim C2 ensures that the top-$p$ candidate sets remain largely stable across layers. Together, these two claims allow us to interpret increases and decreases of entropy as expansion and pruning of the next-token candidate set.
We validate C1 by measuring the correlation between $\Delta H_{i}$ and the change in the number of candidates between layer $i$ and its predecessor. Instead of using Pearson correlation, which assumes a linear relationship, we compute the Spearman rank correlation (Gibbons, 1993). This allows us to test for a monotonic relationship between entropy variations and changes in the number of top-$p$ candidates, without assuming linearity. Table 1 shows a high correlation, confirming the validity of C1. We give further details in Appendix A.1.
Table 1: Spearman rank correlation between layer-wise entropy variations ($\Delta H_{i}$) and changes in the number of next-token candidates across layers, confirming that $\Delta H$ captures meaningful candidate expansion and pruning (Claim C1).
Model Spearman Correlation
Llama-3.2-1B 0.8250
Llama-3.2-3B 0.7365
Gemma-2-2B 0.8804
Gemma-2-9B 0.8540
To validate C2, we measure the percentage of top-$p$ tokens shared by layer $i$ and its predecessor. Figure 1 shows that distributions across layers share many of their candidates, confirming C2.
In summary, the validation of these claims grounds our analysis in a concrete mechanism: it allows us to interpret increments and reductions of entropy as strategies of expansion and pruning of the candidate set, respectively. This justifies the use of $\Delta H$ as an interpretable, low-dimensional view of the token prediction dynamics across layers and for individual generations.
Figure 1: Fraction of top-$p$ ($p=0.6$) next-token candidates shared between a layer and its successor across different models. The consistently high overlap confirms that candidate sets remain stable across layers (Claim C2).
3.3 Three regimes of Rényi entropy
In this subsection, we show that our analysis is not tied to an arbitrary entropy definition, but extends to a broad class of dimensionality reductions for categorical distributions.
Through the main body of the paper, we instantiate our framework using Shannon entropy. More generally, our analysis applies to the broader family of Rényi entropies $H_{\alpha}$, described in Section 2, which subsumes Shannon entropy as the special case $\alpha\rightarrow 1$.
Empirically, we identify an informative regime of $\alpha$ values within which entropy profiles are stable and expressive; importantly, this regime includes the Shannon case $\alpha\rightarrow 1$. In Appendix A.5 and Section 5.3 (Table 3), we show that qualitative and quantitative results remain equivalent within this regime. We therefore adopt Shannon entropy as a robust default for the remainder of the paper, avoiding the need to tune additional hyperparameters.
4 Entropy-Lens
In LLMs, each block operates on the residual stream to map input token sequences to output token distributions. To study their evolution, we introduce Entropy-Lens, a simple, scalable, and model-agnostic framework that summarizes layer-wise predictions through their entropy. As shown in Section 3.2, entropy provides a compact proxy for the number of next-token candidates at each stage of computation.
Conceptually, Entropy-Lens can be seen as a dimensionality reduction of transformer computations, compressing complex layer-by-layer predictions into an entropy profile that provides a compact representation of how the model progressively expands and prunes its space of plausible continuations. In this section, we detail our methodology (further details in Appendix D.1).
Entropy profiles. Formally, let $S=(t_{1},\dots,t_{N})$ denote an input sequence of tokens $t_{j}$ and let $x_{j}^{i}$ be the residual-stream activation of token $t_{j}$ after layer $i$. Similarly to nostalgebraist (2020), decoding this activation through the model’s output head and applying a softmax yields an intermediate probability distribution
$y_{j}^{i}=W(x_{j}^{i}),\quad W:=\operatorname{softmax}\circ D$ (3)
where $D$ is the decoder matrix. We then compute the entropy of this distribution $H(y_{j}^{i})$ to produce a single scalar per layer and token $h_{j}=(H(y_{j}^{1}),\dots,H(y_{j}^{L}))$.
Repeating this computation across layers yields a sequence of entropy values which we call entropy profile. These profiles compress high-dimensional and unordered token prediction dynamics, providing a low-dimensional, layer-wise information-theoretic representation of the model’s computation. Crucially, this construction does not rely on gradients, fine-tuning, or auxiliary probes (Tang et al., 2024; Zhao et al., 2025; Zhang et al., 2026), and can be applied to frozen, off-the-shelf models, in a model agnostic way.
Aggregated entropy profiles. While the entropy profile tracks the candidate pool for a single token, individual predictions are highly variable (Holtzman et al., 2020; Lin et al., 2024). Therefore, model behavior is also analyzed at the level of prompts or generations consisting of multiple tokens. To obtain a stable and informative representation at this level, entropy profiles must be aggregated across tokens.
Specifically, given a prompt, we generate $T$ consecutive tokens autoregressively and extract the profile for each step, yielding a collection of entropy profiles $\{h_{1},\dots,h_{T}\}$. These profiles can be combined using simple aggregation operators to form a single representation of the model’s computation for the entire generation.
In our experiments, we primarily use concatenation across tokens, resulting in a matrix-valued representation that preserves the layer-wise structure of entropy evolution for each generated token. Alternative aggregation strategies, such as averaging across tokens, are also supported. However, we focus on concatenation as it retains maximal temporal information while remaining straightforward to interpret.
Figure 2: Overview of the Entropy-Lens pipeline. Left: intermediate residual-stream activations at each layer are projected into the output space via logit-lens, yielding layer-wise next-token distributions whose entropy is computed. Right: the resulting per-token entropy profiles are aggregated across generated tokens (e.g., by concatenation or averaging) to produce a representation that captures how the model expands and prunes its space of plausible continuations across depth and can optionally be fed to a classifier to assess profile distinctiveness.
In the following sections, we use Entropy-Lens to show that entropy profiles make visible systematic patterns in how transformer models expand and contract their candidate pool across depth. These token prediction dynamics exhibit consistent structure within model families (Section 5.1), and vary across tasks (Section 5.2) and output formats (Section 5.3). Furthermore, the expansion and pruning phases prove to have different relative importance for downstream task performance (Section 5.4).
5 Exploring prediction dynamics in LLMs
The entropy profiles introduced in Section 4 provide a compact view of how intermediate next-token predictions evolve across depth, highlighting phases of expansion and pruning of the candidate set. In this section, we use Entropy-Lens to study the token prediction dynamics in LLMs. We show that different families, tasks and output formats are associated with distinct patterns of expansion and pruning of possible next-tokens.
To analyze the differences in entropy profiles, we employ a k-nearest neighbors (kNN) classifier operating directly on the aggregated entropy representations defined in Section 4. The kNN is not used here as a proposed predictive model, but as a diagnostic tool: because a kNN relies entirely on distances between the inputs, its performance provides a direct indication of the distinctness of entropy profiles.
Across all experiments in this section, entropy profiles are extracted from frozen models. Unless otherwise stated, we evaluate classification performance using standard cross-validation protocols and report area under the ROC curve (AUC) scores. Additional implementation details and experiments are provided in Appendix B and A.2, respectively.
In the next subsections, we aim to answer the following research questions (RQs):
  1. RQ1
    Do different models combine expansion and pruning strategies differently in their token prediction dynamics? Do similarities in these dynamics correspond to structural similarities across models?
  2. RQ2
    Do these dynamics change depending on the task?
  3. RQ3
    Do these dynamics change depending on the output format?
  4. RQ4
    Given a particular model, what is the relative importance of one strategy over the other?
5.1 Family specific dynamics (RQ1)
We first use Entropy-Lens to analyze token prediction dynamics across models, assessing whether observed similarities align with structural similarities.
Entropy profiles exhibit family-level patterns. We compare entropy profiles across 12 decoder-only language models (GPT, Gemma, Llama, Qwen) spanning a wide range of parameter counts (100M to 9B) and depths. Unless otherwise stated, we use a blank prompt and average over multiple generated tokens to obtain a stable profile for each model (details in Appendix B.1).
Figure 3 reveals a striking pattern: entropy profiles cluster by model family rather than by size. In a low-dimensional projection, models from the same family form coherent groups, while distinct families remain well-separated. This indicates that prediction dynamics are not idiosyncratic to individual checkpoints, but reflect a shared computational structure induced by architectural and training choices.
Qualitatively, different families exhibit distinct entropy regimes, as exemplified in Figure 4. GPT models start with high entropy and gradually sharpen their predictions, while Llamas display low-entropy initial stages followed by extended high-entropy plateaus and a final refinement. These differences suggest that model families combine expansion and pruning strategies in systematically different ways.
Depth-rescaling invariance and coarse-to-fine behavior. A remarkable regularity emerges when comparing models of different scales in the same family. While absolute depth varies substantially, entropy profiles exhibit a striking alignment when mapped to a relative layer index (Figure 4). This suggests that model scale primarily dictates the resolution of the computation rather than its qualitative form. From this perspective, smaller models (less than 4B parameters) appear to implement a coarser discretization of the token prediction dynamics found in their larger counterparts: additional layers serve to refine the signal without fundamentally altering the underlying sequence of expansion and pruning.
Overall, these findings establish entropy profiles as readable fingerprints of model-specific token prediction dynamics.
Figure 3: t-SNE projection of aggregated entropy profiles obtained from a blank prompt for different model families. Profiles cluster by family rather than by model size, indicating family-specific token prediction dynamics. Point size scales with number of parameters.
Figure 4: Average entropy profiles over 32 generated tokens for different model families. When depth is normalized, models within the same family exhibit aligned entropy trajectories, revealing family-specific coarse-to-fine token prediction dynamics.
5.2 Task specific dynamics (RQ2)
We now use our framework to assess whether different task types have characteristic token prediction dynamics. We consider three task types that differ in the nature of computation they require: generative (continue a story), syntactic (count the number of words), and semantic (extract the main idea/subject).
Dataset and prompts. We build prompts using the TinyStories dataset (Eldan and Li, 2023). For each task type, we construct prompts by combining a task instruction with a story, using three templates designed to reduce surface-level confounds: Base (task prompt + story), Reversed (story + task prompt), and Scrambled (either task prompt + scrambled story or scrambled story + task prompt), where a scrambled story is obtained by randomly permuting the words in a story. The scrambled condition acts as a control to verify that the entropy signature is driven by the task’s semantic and syntactic requirements rather than simple lexical patterns. Additionally, to further control for prompt phrasing, we use two semantically equivalent variants for each task instruction,
as detailed in Appendix B.2.
Protocol. We generate 800 prompts per task type (2400 total), balanced across templates. For each prompt, we extract aggregated entropy profiles as in Section 4 and evaluate task-type separability using a kNN classifier as a diagnostic probe. We report one-vs-rest ROC-AUC under 10-fold cross-validation.
Results. Across all evaluated models, Table 2 shows that the kNN achieves high AUC, indicating that entropy profiles, and thus token prediction dynamics, present task-relevant structure. Appendix B.2 gives additional details and visualizations, while Appendix A.3 shows the benefits of tracking the entire entropy profile.
Table 2: Task-type classification results on TinyStories using kNN over aggregated entropy profiles. High ROC-AUC scores indicate that entropy and token prediction dynamics exhibit task-specific characteristics.
Model Size k-NN AUC (3 classes)
Gemma-2-it 2.1B 97.66 $\pm$ 0.47
Gemma-2-it 8.9B 98.38 $\pm$ 0.50
Llama-3.2-it 1B 94.94 $\pm$ 0.79
Llama-3.2-it 3B 94.77 $\pm$ 0.93
Llama-3-it 8B 96.10 $\pm$ 0.67
Phi-3 3.6B 97.07 $\pm$ 0.87
5.3 Output format specific dynamics (RQ3)
Similarly to Section 5.2, we now use Entropy-Lens to assess whether different formats of generated text have characteristic token prediction dynamics, independently of topical content. We construct a topic-format dataset by prompting models to generate short texts across multiple topics while enforcing one of three different formats among poem, scientific piece and chat log.
Dataset construction. Each prompt is formed by pairing a format instruction with a topic, as described in Appendix B.4. For each generation, we extract aggregated entropy profiles over the generated text according to Section 4.
Protocol and results. We evaluate format separability using the same kNN diagnostic setup as above and report ROC-AUC under 10-fold cross-validation. Entropy profiles exhibit format-specific characteristics, enabling reliable discrimination between output formats across models (Table 3). We further find that the results remain stable when replacing Shannon entropy with Rényi entropies across a broad range of $\alpha$ values. To visualize similarities, we project aggregated profiles via PCA. The clear clustering in Figure 5 demonstrates linear separability, confirming that format-specific computation leaves a distinctive signature in the token prediction dynamics.
Table 3: Format classification results on the topic–format dataset using kNN over aggregated entropy profiles, computed with different Rényi $\alpha$ values (with Shannon entropy corresponding to $\alpha=1$). Performance is stable across $\alpha$, and high ROC-AUC scores indicate that different output formats are associated with characteristic entropy and token prediction dynamics.
Model $\boldsymbol{\alpha}$ k-NN AUC (3 classes)
0.5 97.3 $\pm$ 1.6
Gemma-2-2B-it 1.0 98.7 $\pm$ 1.1
5.0 98.4 $\pm$ 1.7
0.5 97.8 $\pm$ 1.6
Llama-3.2-1B-it 1.0 97.8 $\pm$ 2.4
5.0 96.6 $\pm$ 2.6
Overall, we have shown that Entropy-Lens proves to be a meaningful tool for the dimensionality reduction of the categorical distributions produced at each transformer layer. Moreover, it enables the analysis of how models expand and prune the set of candidate continuations, revealing shared token prediction dynamics within architectural families and distinct dynamics across tasks and output formats. In the next section, we study how forcing a model to alter its dynamics affects its performance.
Figure 5: PCA projection of aggregated entropy profiles extracted from the topic–format dataset. Profiles cluster by output format, indicating that different formats are associated with characteristic entropy and token prediction dynamics.
5.4 Intervening on token prediction dynamics (RQ4)
Figure 6: Change in MMLU accuracy when skipping fractions of transformer blocks selected by maximal entropy increase (expansion, $\max\Delta H$), maximal entropy decrease (pruning, $\min\Delta H$), or at random. Disrupting expansion phases is generally more detrimental to downstream accuracy, with a notable exception in Gemma2-2B-it.
Entropy-Lens naturally decomposes token prediction dynamics into two strategies: candidate expansion and pruning (as discussed in Section 4). In this section, we assess the relative importance for downstream accuracy of these two strategies by selectively disrupting layers associated with maximal entropy increases ($\max\Delta H$) and decreases ($\min\Delta H$), respectively. We conduct this analysis in a multiple-choice question setting using the MMLU benchmark (Hendrycks et al., 2021). The choice of this dataset is due to several convenient qualities: its strong fact-based components retain good performance on base models, and it possesses a mild reasoning component, essential in a 1-shot setting. We consider six LLMs, including both base and instruction-tuned (it) variants: Gemma2-2B-base, Gemma2-9B-base, Gemma2-2B-it, Gemma-9B-it, Llama3.2-3B-base and Llama3.2-1B-it.
Setup. For each MMLU question, we evaluate the LLM’s prediction by the highest probability token among {A, B, C, D} and comparing it to ground truth. To probe the functional roles of expansion and pruning, we identify the top-$k$ layers that maximize or minimize $\Delta H$ per question. These layers are skipped during inference to measure the impact on accuracy. We vary $k$ from 10% to 50% of the total layers in 10% increments. We employ random layer skipping as a baseline to control for the performance degradation caused by model truncation. This allows us to determine if our targeted strategies yield impacts beyond simple architectural degradation. Furthermore, a random baseline avoids the biases inherent in sequential or fixed-order deletion strategies. In this setting, skipping layer $i$ is implemented by setting the contribution of the $i$-th transformer block to the residual stream to zero.
Results. Figure 6 shows that expansion ($\max\Delta H$) and pruning strategies ($\min\Delta H$) have different relative impact for downstream performance. Across all Llama variants, we observe a critical dependency: skipping the single layer corresponding to maximal expansion causes performance to collapse to random chance. For Gemma2-2B-Base, Gemma2-9B-Base, and Gemma2-9B-it, we observe a distinct separation between the effects of disrupting expansion versus pruning phases, with the former being more detrimental. In contrast, Gemma2-2B-it exhibits a notable inversion of this dynamic, where pruning phases prove significantly more critical for downstream accuracy than expansion phases. Whether this reversal in functional relevance is a direct byproduct of specific alignment or fine-tuning mechanisms remains an open question for future work.